# GIST: Gauge-Invariant Spectral Transformers for Scalable Graph Neural Operators

### 17 Mar 2026 v1

@@ A.1.1 Stage 1: Spectral Convergence and Green’s Function
A.1.2 Stage 2: Random Projection Error
@@
Let $R\in\mathbb{R}^{r\times N}$ be a random projection with $r=\mathcal{O}(\log(N)/\varepsilon^{2})$ constructed via FastRP. For any vectors $v,w\in\mathbb{R}^{N}$,
$\mathbb{P}\left(\left|\langle Rv,Rw\rangle-\langle v,w\rangle\right|\leq\varepsilon\|v\|\|w\|\right)\geq 1-2e^{-c\varepsilon^{2}r}.$
@@
This follows from the Johnson-Lindenstrauss Lemma (Dasgupta2003). The key properties:
  1. 1.
    Random projections with $r=\mathcal{O}(\log(N)/\varepsilon^{2})$ distort distances by at most a factor $(1\pm\varepsilon)$ with high probability.
  2. 2.
    For inner products, since distances are preserved, we have $\langle Rv,Rw\rangle=\frac{1}{2}(\|Rv\|^{2}+\|Rw\|^{2}-\|Rv-Rw\|^{2})\approx\frac{1}{2}(\|v\|^{2}+\|w\|^{2}-\|v-w\|^{2})=\langle v,w\rangle$.
  3. 3.
    FastRP specifically uses sparse random matrices that maintain these guarantees while enabling efficient computation (Chen2019e).
See Dasgupta2003 and Chen2019e for details on FastRP. ∎
A.1.3 Stage 3: Gauge Invariance
Proposition A.4.
GIST’s learned parameters $\theta$ (projection matrix, transformer weights) do not depend on arbitrary gauge choices (sign flips, rotations) in the spectral decomposition because GIST’s computations depend only on gauge-invariant quantities.
Proof.
GIST’s core attention mechanism is:
$\alpha_{ij}=\text{softmax}\left(\frac{\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle}{\sqrt{r}}\right),\quad\tilde{\phi}_{i}=R\phi_{i},$
where $R$ is the random projection matrix. The key insight is that while $\tilde{\phi}_{i}$ depends on $R$ (the arbitrary gauge choice), the inner products $\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle$ do not (in the limit):
$\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle=\phi_{i}^{\top}(R^{\top}R)\phi_{j}\approx\phi_{i}^{\top}\phi_{j},$
by Johnson-Lindenstrauss. Thus $\alpha_{ij}$ converges to a gauge-invariant quantity (the continuum Green’s function kernel) independent of the choice of $R$.
Since all downstream computations operate on gauge-invariant quantities, the learned parameters $\theta$ do not encode any information about the specific gauge choice in the eigenvector decomposition. ∎
@@ A.1.4 Discretization Mismatch Error Analysis
Combining all three stages, we establish the discretization mismatch error bound stated in Proposition A.1(ii) (Gao2025). For two discretizations $\mathcal{G}_{n}$ and $\mathcal{G}_{n^{\prime}}$ of the same manifold $M$, the attention kernel mismatch is bounded by triangle inequality:
$\left|\langle\tilde{\phi}_{i}^{n},\tilde{\phi}_{j}^{n}\rangle-\langle\tilde{\phi}_{i}^{n^{\prime}},\tilde{\phi}_{j}^{n^{\prime}}\rangle\right|\leq\left|\langle\tilde{\phi}_{i}^{n},\tilde{\phi}_{j}^{n}\rangle-G_{M}(x_{i},x_{j})\right|+\left|G_{M}(x_{i},x_{j})-\langle\tilde{\phi}_{i}^{n^{\prime}},\tilde{\phi}_{j}^{n^{\prime}}\rangle\right|.$
Each term on the right decomposes into spectral convergence error (Stage 1) and random projection error (Stage 2), yielding the total bound $\mathcal{O}(n^{-1/(m+4)})$ where $n$ is the coarser discretization. For typical manifold dimensions, random projection errors decay faster than spectral convergence errors, so the latter dominate the discretization mismatch.
A.2 Pseudo-code
@@
2:Matrix of NN node graph positional embeddings $\mathbf{\Phi}\in\mathbb{R}^{N\times r}$
@@
2:Output sequence $\mathbf{O}\in\mathbb{R}^{N\times d}$ to be applied to graph positional embeddings $\mathbf{\Phi}$
@@
This property is crucial for preserving gauge invariance: since the attention weights are computed as $\langle\varphi(\tilde{\phi}_{i}),\varphi(\tilde{\phi}_{j})\rangle\approx k_{0}(\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j})$ and $\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j}\approx\phi_{i}^{\top}\phi_{j}$ by Johnson-Lindenstrauss (as established in Section A), the resulting attention pattern depends only on gauge-invariant inner products between the spectral embeddings.
@@
A.3 GIST Hyperparameters robustness
In order to study the sensitivity of GIST’s performance to variations in its spectral embedding hyperparameters, we train multiple simplified GIST architectures (two-block Gauge-Invariant Spectral Self-Attention linear transformers) on the Cora benchmark while varying the power iteration parameter $k$ and the embedding dimension $r$ in the FastRP approximation. These parameters directly control the quality of spectral embeddings while balancing computational efficiency.
@@ A.3 GIST Hyperparameters robustness
A.4 GIST Scalability Study
@@
A.5 Multi-Scale Architecture Ablation Study

### 10 Apr 2026 v2

@@ A.1.1 Stage 1: Spectral Convergence and Green’s Function
Corollary A.3 (Kernel convergence for general spectral filters).
Let $f$ be a spectral filter such that $g(\lambda):=f(\lambda)^{2}$ is Lipschitz on the spectrum of $\Delta_{M}$, with Lipschitz constant $L_{g}$. Then the discrete kernel $K_{n}(i,j)=\sum_{l}f(\lambda_{l}^{n})^{2}\,(u_{l}^{n})_{i}(u_{l}^{n})_{j}$ converges to the continuum kernel $K_{\infty}(x_{i},x_{j}):=\sum_{l}f(\mu_{l})^{2}\,\psi_{l}(x_{i})\psi_{l}(x_{j})$ at rate $\mathcal{O}(n^{-1/(m+4)})$.
Proof.
By the spectral convergence of the previous proposition (part a), $|\lambda_{k}^{n}-\mu_{k}|=\mathcal{O}(n^{-1/(m+4)})$ and $\|u_{k}^{n}-\psi_{k}\|_{L^{2}}=\mathcal{O}(n^{-1/(m+4)})$. The Lipschitz condition on $g$ gives $|g(\lambda_{k}^{n})-g(\mu_{k})|\leq L_{g}|\lambda_{k}^{n}-\mu_{k}|=\mathcal{O}(n^{-1/(m+4)})$. Combined with the eigenvector convergence (each term in the sum has both eigenvalue and eigenvector errors at the same rate), this yields $|K_{n}(i,j)-K_{\infty}(x_{i},x_{j})|=\mathcal{O}(n^{-1/(m+4)})$. When $f$ approximates $\lambda\mapsto 1/\sqrt{\lambda}$, $K_{\infty}$ reduces to the Green’s function $G_{M}$. ∎
A.1.2 Stage 2: Spectral Embeddings as Kernel Estimators
We consider the general class of approximate spectral embeddings $\widetilde{\Phi}=f(P)\cdot R\in\mathbb{R}^{N\times r}$, where $P$ is a graph operator, $f$ is a spectral filter (see §3 for definitions), and $R\in\mathbb{R}^{N\times r}$ is a random matrix with $\mathbb{E}[RR^{\top}]=\mathds{1}$. We write $\phi_{i}=f(P)_{i\cdot}^{\top}\in\mathbb{R}^{N}$ for the $i$-th row of $f(P)$ (the “exact” spectral embedding of node $i$) and $\tilde{\phi}_{i}\in\mathbb{R}^{r}$ for the $i$-th row of $\widetilde{\Phi}$ (its random projection); Laplacian eigenmaps equation 2 are the special case $f(\lambda)=\lambda^{-1/2}$, $P=\mathcal{L}$.
Lemma A.4 (Kernel estimator).
For any spectral filter $f$ and random matrix $R\in\mathbb{R}^{N\times r}$ with $\mathbb{E}[RR^{\top}]=\mathds{1}$, the inner products of $\widetilde{\Phi}=f(P)\cdot R$ satisfy
$\mathbb{E}\!\left[\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle\right]=K(i,j):=[f(P)f(P)^{\top}]_{ij}.$
When $P$ is symmetric with eigenvectors $u_{l}$, this reduces to $K(i,j)=\sum_{l}f(\lambda_{l})^{2}\,(u_{l})_{i}(u_{l})_{j}$.
Proof.
$\mathbb{E}[\widetilde{\Phi}\,\widetilde{\Phi}^{\top}]=f(P)\,\mathbb{E}[RR^{\top}]\,f(P)^{\top}=f(P)\cdot\mathds{1}\cdot f(P)^{\top}=f(P)f(P)^{\top}$. Taking the $(i,j)$ entry gives the result. ∎
Proposition A.5 (Exact gauge invariance of $K$).
The kernel $K(i,j)=[f(P)f(P)^{\top}]_{ij}$ is exactly invariant under all gauge transformations of the eigenvector basis: sign flips $u_{l}\to s_{l}u_{l}$, orthogonal rotations within degenerate eigenspaces, and different numerical solver choices.
Proof.
$K(i,j)=[f(P)f(P)^{\top}]_{ij}$ is an entry of the matrix $f(P)f(P)^{\top}$, which is a function of $P$ alone and does not depend on the choice of eigenvector basis. All gauge transformations leave $P$ unchanged, so $K$ is unchanged. ∎
Lemma A.6 (Concentration bound).
Let $\mathbf{a}=f(P)_{i\cdot}^{\top}$ and $\mathbf{b}=f(P)_{j\cdot}^{\top}$ be the $i$-th and $j$-th rows of $f(P)$, and let $R\in\mathbb{R}^{N\times r}$ have i.i.d. sub-Gaussian entries with mean zero and variance $1/r$. Then:
(i) Tail bound. By the Hanson-Wright inequality (Rudelson2013), for any $t>0$:
$\Pr\!\left[\left|\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle-K(i,j)\right|>t\right]\leq 2\exp\!\left(-c\cdot r\cdot\min\!\left(\frac{t^{2}}{\|\mathbf{a}\|^{2}\|\mathbf{b}\|^{2}},\frac{t}{\|\mathbf{a}\|\|\mathbf{b}\|}\right)\right),$
where $c>0$ is an absolute constant depending only on the sub-Gaussian norm of the entries of $R$.
(ii) Scaling of $r$. Fix a target accuracy $\varepsilon>0$ and a failure probability $\delta\in(0,1)$. Setting $t=\varepsilon\|\mathbf{a}\|\|\mathbf{b}\|$ in the bound above, and noting that for $\varepsilon\leq 1$ the minimum is achieved by the quadratic term $t^{2}/(\|\mathbf{a}\|^{2}\|\mathbf{b}\|^{2})=\varepsilon^{2}$, we require
$2\exp(-c\cdot r\cdot\varepsilon^{2})\leq\delta\qquad\Longleftrightarrow\qquad r\geq\frac{\log(2/\delta)}{c\,\varepsilon^{2}}.$
Therefore $r=\mathcal{O}(\varepsilon^{-2}\log(1/\delta))$ random-projection dimensions suffice to guarantee
$\left|\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle-K(i,j)\right|\leq\varepsilon\,\|\mathbf{a}\|\|\mathbf{b}\|$
with probability at least $1-\delta$. Equivalently, for fixed $\delta$, the estimation error is $\mathcal{O}(\|\mathbf{a}\|\|\mathbf{b}\|/\sqrt{r})$.
Proof.
Write $R=[\mathbf{r}_{1}\;\cdots\;\mathbf{r}_{r}]$ where each column $\mathbf{r}_{s}\in\mathbb{R}^{N}$ has i.i.d. sub-Gaussian entries with mean zero and variance $1/r$. Then
$\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle=\sum_{s=1}^{r}(\mathbf{a}^{\top}\mathbf{r}_{s})(\mathbf{b}^{\top}\mathbf{r}_{s})=\mathbf{a}^{\top}\!\left(\sum_{s=1}^{r}\mathbf{r}_{s}\mathbf{r}_{s}^{\top}\right)\!\mathbf{b}=\mathbf{a}^{\top}(RR^{\top})\mathbf{b}.$
Each summand $(\mathbf{a}^{\top}\mathbf{r}_{s})(\mathbf{b}^{\top}\mathbf{r}_{s})$ is independent with mean $\mathbb{E}[\mathbf{a}^{\top}\mathbf{r}_{s}\mathbf{r}_{s}^{\top}\mathbf{b}]=(1/r)\,\mathbf{a}^{\top}\mathbf{b}=(1/r)\,K(i,j)$, so $\mathbb{E}[\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle]=K(i,j)$. The product $(\mathbf{a}^{\top}\mathbf{r}_{s})(\mathbf{b}^{\top}\mathbf{r}_{s})$ can be written as $\mathbf{r}_{s}^{\top}(\mathbf{a}\mathbf{b}^{\top})\mathbf{r}_{s}$, a quadratic form in a sub-Gaussian random vector; summing $r$ such terms and applying the Hanson-Wright inequality (Rudelson2013) to $\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle-K(i,j)=\mathbf{a}^{\top}(RR^{\top}-\mathds{1})\mathbf{b}$ yields the tail bound in part (i). Part (ii) follows by inverting this bound. ∎
A.1.3 Stage 3: Gauge Invariance of GIST
@@
GIST’s learned parameters do not depend on arbitrary gauge choices in the spectral decomposition, because GIST’s attention depends only on the exactly gauge-invariant kernel $K(i,j)$.
@@
GIST computes attention weights via linear attention with a feature map $\varphi$: the effective logits are $e_{ij}=\langle\varphi(\tilde{\phi}_{i}),\varphi(\tilde{\phi}_{j})\rangle$. By Lemma A.4, $\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle$ estimates $K(i,j)=[f(P)f(P)^{\top}]_{ij}$, and by Proposition A.5, $K(i,j)$ is exactly gauge-invariant. When $\varphi=\mathrm{ReLU}$ (as in Algorithm 2), the inner product $\langle\varphi(\tilde{\phi}_{i}),\varphi(\tilde{\phi}_{j})\rangle$ converges to the arc-cosine kernel $k_{0}(\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j})$ as $r\to\infty$ (Cho2009), which depends only on the inner product $\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j}\to K(i,j)$ and is therefore gauge-invariant. All downstream computations operate on these gauge-invariant logits and on node features $x_{i}$ (which are gauge-independent), so the full architecture preserves gauge invariance. ∎
@@ A.1.4 Discretization Mismatch Error Analysis
Combining all three stages, for two discretizations $\mathcal{G}_{n}$ and $\mathcal{G}_{n^{\prime}}$ of the same manifold $M$, the attention kernel mismatch is bounded by the triangle inequality:
[…]
where $K_{n}(i,j)=[f(P_{n})f(P_{n})^{\top}]_{ij}$ denotes the kernel on discretization $\mathcal{G}_{n}$. The middle term follows from Corollary A.3: both $K_{n}$ and $K_{n^{\prime}}$ converge to the same continuum kernel $K_{\infty}(x_{i},x_{j})=\sum_{l}f(\mu_{l})^{2}\psi_{l}(x_{i})\psi_{l}(x_{j})$ at rate $\mathcal{O}(n^{-1/(m+4)})$, so by the triangle inequality $|K_{n}(i,j)-K_{n^{\prime}}(i,j)|=\mathcal{O}(n^{-1/(m+4)})$. The total bound is therefore
$\left|\langle\tilde{\phi}_{i}^{n},\tilde{\phi}_{j}^{n}\rangle-\langle\tilde{\phi}_{i}^{n^{\prime}},\tilde{\phi}_{j}^{n^{\prime}}\rangle\right|=\mathcal{O}\!\left(\frac{1}{\sqrt{r}}\right)+\mathcal{O}\!\left(n^{-1/(m+4)}\right),$
with the two error sources being independent. Choosing $r=\mathcal{O}(\varepsilon^{-2})$ and $n=\mathcal{O}(\varepsilon^{-(m+4)})$ achieves total error $\varepsilon$.
Appendix B Algorithms
@@
2:Matrix of $N$ node graph positional embeddings $\mathbf{\Phi}\in\mathbb{R}^{N\times r}$
@@
2:Output sequence $\mathbf{O}\in\mathbb{R}^{N\times{\color[rgb]{1,0,0}r}}$ to be applied to graph positional embeddings $\mathbf{\Phi}$
@@
This property is crucial for preserving gauge invariance: the attention weights $\langle\varphi(\tilde{\phi}_{i}),\varphi(\tilde{\phi}_{j})\rangle\approx k_{0}(\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j})$ depend only on the inner product $\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j}$, which by Lemma A.4 estimates the exactly gauge-invariant kernel $K(i,j)=[f(P)f(P)^{\top}]_{ij}$. The resulting attention pattern therefore depends only on gauge-invariant quantities.
@@
Appendix C Additional Empirical Studies
C.1 GIST Hyperparameters robustness
In order to study the sensitivity of GIST’s performance to variations in its spectral embedding hyperparameters, we train multiple simplified GIST architectures (two-block Gauge-Invariant Spectral Self-Attention linear transformers) on the Cora benchmark while varying the power iteration parameter $k$ and the embedding dimension $r$ of the spectral embedding. These parameters directly control the quality of spectral embeddings while balancing computational efficiency.
@@ C.1 GIST Hyperparameters robustness
C.2 GIST Scalability Study
@@
C.3 Multi-Scale Architecture Ablation Study
@@
C.4 Pronunciation
GIST is pronounced with a soft G, like gist (the English word for the central point or essence of something), not with a hard G as in gift. Although insisting on a hard G won’t necessarily be considered wrong. After all, that is just a different gauge choice.